# Anomalous enhancement of Néel order in the $S=1/2$ square lattice Heisenberg model under fictitious magnetic field

### v1

@@ Abstract
Fictitious magnetic fields can be introduced in quantum magnets by the strain engineering and the Aharonov-Casher effect. Here, we study impacts of a uniform fictitious magnetic field and corresponding Landau quantization on the Néel order in the $S=1/2$ Heisenberg model on the square lattice as a prototypical system of quantum magnets. We first analyze the system by using the spin-wave approximation. It is found that the staggered magnetization is enhanced by the fictitious magnetic field and it shows a simple scaling behavior with respect to the magnetic length, in contrast to a naive expectation based on the Dzyaloshinskii-Moriya interaction. We also perform density matrix renormalization group calculations to fully take quantum effects into account and obtain quantitatively more accurate results. The enhancement is found to be anomalously strong and it shows a non-trivial scaling behavior or a fluctuation induced discontinuity. These results are beyond the well known understanding of correlated systems under magnetic fields called the magnetic catalysis.
@@ I Introduction
A magnetic field has significant impacts on physical properties not only through the Zeeman effect but also through the orbital effect. The orbital effect is related to the Landau quantization of a band structure and it has been studied mainly for itinerant electron systems. However, a fictitious magnetic field and corresponding Landau quantization can be introduced in electrically charge-neutral insulators. For example, a fictitious magnetic field can be introduced by strain. The strain-induced fictitious magnetic field was first discussed for graphene, where triaxial strain can effectively lead to a non-uniform magnetic field [1, 2]. Similar mechanisms work also for electrically charge neutral quasi-particles such as emergent Majorana fermions and bosonic magnons in quantum spin systems on the honeycomb lattice, because strain generally leads to modulation of hopping integrals and electromagnetic coupling is not necessary [3, 4, 5, 6, 7, 8]. The strain-induced fictitious magnetic fields were generalized to other lattices such as Kagome, $\alpha$-$T_{3}$, and square lattices [9, 10, 11, 12].
@@
The other promising approach for realization of a fictitious magnetic field in a quantum magnet is the Aharonov-Casher effect [13, 14]. In a system with the spin-orbit interaction, an electric field adds to momentum of electrons and thus can behave as an SU(2) gauge field in the spin space, which modifies magnetic interactions and leads to Dzyaloshinskii-Moriya like interactions. Such an effect has been evaluated theoretically [15, 16] and examined experimentally for magnets [17, 18]. The fundamental difference between the strain engineering and the Aharonov-Casher effect is that a spatially uniform fictitious magnetic field can be realized and spin rotation symmetry is reduced in the latter. For a one-dimensional spin chain, the Aharonov-Casher effect with a uniform electric field is simply implemented by the twisted boundary condition and field dependence of the spin current in the Heisenberg model was studied with using the Bethe ansatz [19]. In two dimensions, a uniform electric field generates a constant Aharonov-Casher phase in the magnetic interaction, while an electric field gradient can induce a fictitious magnetic flux which leads to Landau quantization of magnons [20, 21, 22]. Theoretically, within the spin-wave approximation neglecting magnon interactions, it was shown that magnons under a fictitious magnetic field can have topologically non-trivial bands and exhibit various Hall effects. It was also proposed that a fictitious magnetic field can give rise to the spin Nernst effect of magnons based on the spin-wave approximation [23]. The basic physics behind these interesting transport phenomena is controlling magnon bands, which is called magnonics [24]. In this context, the fictitious magnetic field induced by the Aharonov-Casher effect can be a useful tool for engineering magnons.
@@
However, despite the extensive studies of quantum spin systems with fictitious magnetic fields, their impacts on magnetic orders have not been well understood. Especially, it is a non-trivial problem whether or not the magnon interactions and quantum fluctuations often ignored within the spin-wave approximation are indeed safely negligible. It was shown that, in the $S=1/2$ Heisenberg model on the honeycomb lattice under a strain induced non-uniform field, the antiferromagnetic order is weakened even around the center of the system and suppression in the quantum Monte Carlo calculations is much stronger than that in the spin-wave approximation, implying the importance of quantum fluctuations [8]. On the other hand, to our knowledge, magnetism under a uniform fictitious magnetic field has not been theoretically explored and the fate of the magnetism is not known. In contrast to strain-induced fields, a fictitious magnetic field by the Aharonov-Casher effect reduces the spin rotation symmetry and thus a magnetic order could show different behaviors from those under strain. For an electric field which is uniform within a plane, the corresponding fictitious vector potential can be approximated by a Dzyaloshinskii-Moriya interaction and it will suppress the Néel order. In presence of an electric field gradient which gives a fictitious magnetic flux, however, validity of such an approximation is subtle and enhancement of the Néel order might be possible. Furthermore, a change of the magnetic order could be strengthened by magnon interactions.
@@
In this study, we discuss effects of uniform fictitious magnetic fields in the $S=1/2$ Heisenberg model on the square lattice as a prototypical model of quantum spin systems. We first analyze the model using the spin-wave approximation, and next perform density-matrix-renormalization-group (DMRG) calculations to fully take quantum effects into account and obtain quantitatively more accurate results [25, 26, 27]. It is found that the Néel order is enhanced by the fictitious magnetic field in contrast to the above-mentioned naive expectation based on the uniform Dzyaloshinskii-Moriya interaction. Furthermore, the DMRG calculation shows much stronger enhancement of the Néel order, which indicates importance of the quantum fluctuations neglected in the spin-wave approximation. We point out that this enhancement could be regarded as a bosonic variant of the magnetic catalysis which has been extensively studied mainly for Dirac fermion systems [28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. However, the anomalous enhancement has not been seen in any fermion systems and is beyond the standard magnetic catalysis. The fictitious magnetic field in quantum magnets could provide an interesting platform for the interplay of Landau quantization and interactions.
@@ II Model and Calculation Method
We study the $S=1/2$ antiferromagnetic Heisenberg model on the square lattice (system size $L_{x}\times L_{y}$) under a fictitious magnetic flux $\phi$ as shown in Fig. 1. The magnetic field is uniform over the entire system, $\sum_{ij\in p}\theta_{ij}=2\pi\phi$ for every plaquette $p$. Note that the lattice constant is the length unit in our model and there is no distinction between a magnetic field and a magnetic flux per plaquette. The fictitious magnetic field could be realized by an electric field gradient within the plane through the Aharonov-Casher effect or alternatively by an inhomogeneous Dyzaloshinskii-Moriya interaction [20, 21, 22, 23, 24]. The Hamiltonian is
$\mathcal{H}=\sum_{\expectationvalue{i,j}}\left[\frac{1}{2}\left(e^{i{\theta}_{ij}}S^{+}_{i}S^{-}_{j}+e^{-i{\theta}_{ij}}S^{-}_{i}S^{+}_{j}\right)+S^{z}_{i}S^{z}_{j}\right],$ (1)
where $J>0$ is an antiferromagnetic interaction and $\expectationvalue{i,j}$ is a pair of nearest-neighbor sites. The $S_{j}^{\pm}$ operators are $S_{j}^{\pm}=S_{j}^{x}\pm iS_{j}^{y}$. The present magnetic field is a fictitious field and the Hamiltonian does not contain a Zeeman term. We introduce the fictitious vector potential $\theta_{ij}$ in the Landau gauge as
@@
[…]
@@
In this study, we first apply the spin-wave approximation which is a standard method to describe low-energy properties of quantum magnets. The Hamiltonian is approximated within quadratic terms in the boson operators and interactions between the magnons are neglected (see Appendix A). In the spin-wave approximation, both of the torus and the cylinder geometries are used depending on physical quantities considered and ranges of the fictitious magnetic flux. Next, we perform the DMRG calculation in the cylinder geometry to obtain more accurate results. The quantum fluctuations arising from the magnon interaction neglected in the spin-wave approximation are taken into account in the DMRG calculation. The aspect ratio is chosen as $(L_{x},L_{y})=(2L,L)$ with the system size $L=6,8,10$. The bond dimension is $\chi=800\sim 3200$ and the truncation error is $\epsilon<10^{-5}$. [26, 25, 27].
@@ III Calculation Results
In this section, we show the calculation results of the spin-wave approximation and the DMRG at zero temperature. They give qualitatively consistent results, but there is significant quantitative difference. To have an insight into the results of DMRG calculation , we give a brief discussion based on the magnetic catalysis.
@@ III.1 Spin-wave approximation
In the spin-wave approximation, we suppose that the classical ground state at $\phi=0$ is the Néel ordered state in the spin $z$-axis. The staggered magnetic moment $m_{s}$ will increase if the ${\mathbb{Z}}_{2}$ symmetry breaking state becomes more stable, while $m_{s}$ would decrease if the U(1) symmetry breaking takes place. In the latter case, the $z$-axis Néel ansatz is inappropriate and the spin-wave approximation based on this ansatz state would not work. More precisely, the energy spectrum of magnons will have imaginary components when the spin-wave approximation becomes invalid. A failure of the spin-wave approximation implies instability of the Néel ordered ansatz and a different order may be realized in the ground state.
We first calculate energy spectrum of the magnons and investigate how the magnon excitations depend on the fictitious magnetic flux $\phi$.
@@
Figure 2 shows the energy spectrum calculated by the spin-wave approximation for various magnetic fluxes $0\leq\phi\leq 1$. The magnon spectrum shows a Hofstadter butterfly as in a free fermion system, as demonstrated in the previous studies [38, 22, 23]. Note that the calculated magnon energy is always real (non-imaginary) for all $\phi$, which implies that the $z$-axis Néel ground state is stable under the fictitious magnetic field. If the $z$-axis Néel state were unstable under some fluxes, the magnon energy would take complex values. One can see that the magnon excitation acquires an energy gap under the magnetic field $\phi\neq 0$. The magnon excitation gap $\Delta E$ (the smallest magnon excitation energy) behaves as $\Delta E\propto\sqrt{\phi}$ for small magnetic fluxes $0<\phi\lesssim 0.01$ as shown in Figs. 3(a) and (b), which is consistent with the long wavelength behavior described by the Klein-Gordon equation [23].
Figure 3: Magnon energy gap for small values of the fictitious magnetic field (a) in the linear scale and (b) in the logarithmic scale. The system size is $L_{x}=L_{y}=100$ sites with the cylinder geometry. The dashed curve in (b) is $\Delta E\sim\sqrt{\phi}$. Small deviation from the exact $\Delta E\sim\sqrt{\phi}$ behavior is considered to be a finite size effect.
Figure 4: (a) Néel order parameter $m_{s}$ calculated by the spin-wave approximation in the cylinder geometry for the system size $L_{x}=100,L_{y}=50$. The inset is the difference $\Delta m_{s}$ for small fluxes $0\leqq\phi\lesssim 0.01$. (b) Logarithmic scale plot of $\Delta m_{s}$. The dashed line represents $\Delta m_{s}\propto\sqrt{\phi}$.
The $\Delta E\propto\sqrt{\phi}$ dependence can be understood by a simple dimensional analysis of the linear spectra in the momentum space $\Delta E\propto k$ at $\phi=0$. Indeed, $\Delta E\propto k\propto 1/L$ for the system size $L$ and one can just replace $L$ with the magnetic length $l_{B}=1/\sqrt{\phi}$ to obtain $\Delta E\propto 1/l_{B}=\sqrt{\phi}$. Note that the long-wave length limit of the magnon spectrum is effectively described by the Klein-Gordon equation and there is no zero mode ($\Delta E=0$) in sharp contrast to Dirac fermions at charge neutrality. The non-zero magnon gap implies that any excitations from the ground state are gapped since the magnons are elementary excitations in the Heisenberg model. The correlation length will be given by $\xi\sim v/\Delta E\propto l_{B}$ where $v$ is the speed of the magnon excitation, and the system is governed by this length scale.
@@
where we use use the cylinder geometry since it allows small fluxes. $N$ is the number of sites in the summation and it is restricted to the $N=L\times L$ sites in the middle of the system for the cylinder geometry with $(L_{x},L_{y})=(2L,L)$ to suppress contributions from the open boundary. Note that the spin-wave approximation works quite well for the Heisenberg model at $\phi=0$ and it gives $m_{s}\simeq 0.303$ in the thermodynamic limit which is very close to $m_{s}\simeq 0.306\sim 0.309$ obtained by the extensive numerical calculations [39, 40, 41, 42, 43, 44, 45]. Figure 4(a) shows the Néel order parameter $m_{s}$ under the fictitious magnetic flux $\phi$. We find that the Néel order parameter is enhanced from the zero flux value $m_{s}(0)\simeq 0.303$ and it reaches $m_{s}\sim 0.41$ at large fluxes. Once the enhanced $m_{s}$ has been obtained, it can be naturally understood from the gapped magnon spectrum in analogy to an Ising-like system at $\phi=0$ where magnons are gapped and correspondingly $m_{s}$ is large. Since zero-point motion of the magnons leads to the suppression of $m_{s}$ from the classical value $m_{\rm max}=0.5$ within the spin-wave approximation, a magnon gap would generally imply increase of $m_{s}$. At the same time, however, the enhanced Néel order may be non-trivial, once one notices that the Hamiltonian (1) partly includes a non-uniform Dzyaloshinskii-Moriya interaction in the spin $xy$-plane when $e^{i\theta_{ij}}$ is simply Taylor expanded up to the lowest order in $\phi$. A uniform $xy$-Dyzloshinskii-Moriya interaction is known to destabilize the $z$-axis Néel order, and one might naively expect that the non-uniform Dyzloshinskii-Moriya interaction has a similar influence. The enhancement of the Néel order in our calculation suggests that it is important to fully take the Landau quantization into account beyond the naive Taylor expansion.

### v2

# Anomalous enhancement of Néel order in the $S=\frac{1}{2}$ square lattice Heisenberg model under fictitious magnetic field

@@ Abstract
Fictitious magnetic fields can be introduced in quantum magnets by strain engineering and the Aharonov-Casher effect. Here, we study impacts of a uniform fictitious magnetic field and corresponding Landau quantization on the Néel order in the $S=\frac{1}{2}$ Heisenberg model on the square lattice as a prototypical system of quantum magnets. We first analyze the system by using the spin-wave approximation. It is found that the staggered magnetization is enhanced by the fictitious magnetic field and it shows a simple scaling behavior with respect to the magnetic length, in contrast with a naive expectation based on the Dzyaloshinskii-Moriya interaction. We also perform density matrix renormalization group calculations to fully consider quantum effects and obtain quantitatively more accurate results. The enhancement is found to be anomalously strong and it shows a non trivial scaling behavior or a fluctuation induced discontinuity. These results are beyond the well-known understanding of correlated systems under magnetic fields called the magnetic catalysis.
@@ I Introduction
A magnetic field has significant impacts on physical properties not only through the Zeeman effect but also through the orbital effect. The orbital effect is related to the Landau quantization of a band structure and it has been studied mainly for itinerant electron systems. However, a fictitious magnetic field and corresponding Landau quantization can be introduced in electrically charge-neutral insulators. For example, a fictitious magnetic field can be introduced by strain. The strain-induced fictitious magnetic field was discussed for graphene, where triaxial strain can effectively lead to a non-uniform magnetic field [1, 2]. Similar mechanisms also work for electrically charge-neutral quasi-particles such as emergent Majorana fermions and bosonic magnons in quantum spin systems on the honeycomb lattice, because strain generally leads to modulation of hopping integrals, and electromagnetic coupling is not necessary [3, 4, 5, 6, 7]. The strain-induced fictitious magnetic fields were generalized to other lattices such as kagome, $\alpha$-$T_{3}$, and square lattices [8, 9, 10, 11].
@@
The other promising approach for realization of a fictitious magnetic field in a quantum magnet is the Aharonov-Casher effect [12, 13]. In a system with a spin-orbit interaction, an electric field adds to momentum of electrons and thus can behave as an SU(2) gauge field in the spin space, which modifies magnetic interactions and leads to Dzyaloshinskii-Moriya like interactions. Such an effect has been evaluated theoretically [14, 15] and examined experimentally for magnets [16, 17]. The fundamental difference between the strain engineering and the Aharonov-Casher effect is that a spatially uniform fictitious magnetic field can be realized and spin-rotation symmetry is reduced in the latter. For a one-dimensional spin chain, the Aharonov-Casher effect with a uniform electric field is simply implemented by the twisted boundary condition, and field dependence of the spin current in the Heisenberg model was studied using the Bethe ansatz [18]. In two dimensions, a uniform electric field generates a constant Aharonov-Casher phase in the magnetic interaction, while an electric field gradient can induce a fictitious magnetic flux which leads to Landau quantization of magnons [19, 20, 21]. Theoretically, within the spin-wave approximation neglecting magnon interactions, it was shown that magnons under a fictitious magnetic field can have topologically non-trivial bands and exhibit various Hall effects. It was also proposed that a fictitious magnetic field can give rise to the spin Nernst effect of magnons based on the spin-wave approximation [22]. The basic physics behind these interesting transport phenomena is controlling magnon bands, which is called magnonics [23]. In this context, the fictitious magnetic field induced by the Aharonov-Casher effect can be a useful tool for engineering magnons.
@@
However, despite extensive studies of quantum spin systems with fictitious magnetic fields, their impacts on magnetic orders have not been well understood. It is a non trivial problem whether the magnon interactions and quantum fluctuations often ignored within the spin-wave approximation are indeed safely negligible. It was shown that, in the $S=\frac{1}{2}$ Heisenberg model on the honeycomb lattice under a strain-induced non uniform field, the antiferromagnetic order is weakened near the corners of the system and suppression in the quantum Monte Carlo calculations is much stronger than that in the spin-wave approximation, implying the importance of quantum fluctuations [7]. On the other hand, to our knowledge, magnetism under a uniform fictitious magnetic field has not been theoretically explored, and the fate of the magnetism is not known. In contrast with strain-induced fields, a fictitious magnetic field by the Aharonov-Casher effect reduces the spin rotation symmetry and thus a magnetic order could show different behaviors from those under strain. For an electric field which is uniform within a plane, the corresponding fictitious vector potential can be approximated by a Dzyaloshinskii-Moriya interaction, and it will suppress the Néel order. In the presence of an electric field gradient which gives a fictitious magnetic flux, however, validity of such an approximation is subtle and enhancement of the Néel order might be possible. Furthermore, a change of the magnetic order could be strengthened by magnon interactions.
@@
In this paper, we discuss effects of uniform fictitious magnetic fields in the $S=\frac{1}{2}$ Heisenberg model on the square lattice as a prototypical model of quantum spin systems. We first analyze the model using the spin-wave approximation, and next perform density-matrix-renormalization-group (DMRG) calculations to fully take consider quantum effects into account and obtain quantitatively more accurate results [24, 25, 26]. It is found that the Néel order is enhanced by the fictitious magnetic field in contrast with the above-mentioned naive expectation based on the uniform Dzyaloshinskii-Moriya interaction. Furthermore, the DMRG calculation shows much stronger enhancement of the Néel order, which indicates the importance of the quantum fluctuations neglected in the spin-wave approximation. We point out that this enhancement could be regarded as a bosonic variant of the magnetic catalysis, which has been extensively studied , mainly for Dirac fermion systems [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. However, the anomalous enhancement has not been seen in any fermion systems and is beyond the standard magnetic catalysis. The fictitious magnetic field in quantum magnets could provide an interesting platform for the interplay of Landau quantization and interactions.
@@ II Model and Calculation Method
We study the $S=\frac{1}{2}$ antiferromagnetic Heisenberg model on a square lattice (system size $L_{x}\times L_{y}$) under a fictitious magnetic flux $\phi$, as shown in Fig. 1. Although the fictitious magnetic field in the Aharonov-Casher effect is generally described by an SU(2) gauge field, we focus on the U(1) part in our analysis. The magnetic field is uniform over the entire system, $\sum_{ij\in p}\theta_{ij}=2\pi\phi$ for every plaquette $p$. Similar to the Aharonov-Bohm effect, the spin acquires a phase $2\pi\phi$ when it goes around a plaquette. Note that the lattice constant is the length unit in our model, and there is no distinction between a magnetic field and a magnetic flux per plaquette, nor between the vector potential and the phase factor. The fictitious magnetic field could be realized by an electric field gradient within the plane through the Aharonov-Casher effect or alternatively by an inhomogeneous Dyzaloshinskii-Moriya interaction [19, 20, 21, 22, 23]. The Hamiltonian is
$\mathcal{H}=J\sum_{\expectationvalue{i,j}}\left[\frac{1}{2}\left(e^{i{\theta}_{ij}}S^{+}_{i}S^{-}_{j}+e^{-i{\theta}_{ij}}S^{-}_{i}S^{+}_{j}\right)+S^{z}_{i}S^{z}_{j}\right],$ (1)
where $J>0$ is an antiferromagnetic interaction, and $\expectationvalue{i,j}$ is a pair of nearest-neighbor sites. The $S_{j}^{\pm}$ operators are $S_{j}^{\pm}=S_{j}^{x}\pm iS_{j}^{y}$. The present magnetic field is a fictitious field and the Hamiltonian does not contain a Zeeman term. We introduce the fictitious vector potential $\theta_{ij}$ in the Landau gauge as
@@
where the site position is denoted by $\bm{r}_{j}=(x_{j},y_{j})$ with $x_{j}=0,1,\dots,L_{x}-1$ and $y_{j}=0,1,\dots,L_{y}-1$. The Heisenberg model with the vector potential has U(1) global spin-rotation symmetry along the spin-$z$axis, and the spin-$\pi$ rotation symmetry along the-$x$ axis combined with the complex conjugation, $z\to z^{*},S_{j}^{+}\to S_{j}^{-},S_{j}^{z}\to-S_{j}^{z}$, where $z$ is a complex constant. Note that, in the standard notation of the spin operators with the Pauli matrices, the $y$-component transforms as $S_{j}^{y}\to-S_{j}^{y}$ under the complex conjugation. Thus, SU(2) spin-rotation symmetry at $\phi=0$ is reduced to ${\mathrm{U}}(1)\times{\mathbb{Z}}_{2}$ symmetry in the presence of $\phi\neq 0$ (the $\mathbb{Z}_{2}$ symmetry is anti-unitary). The combined-$\pi$ rotation symmetry is spontaneously broken in the $z$-axis Néel ordered state. In this paper, we use two system geometries, namely a torus and a cylinder, where there are no boundary effects in the former while the latter is convenient for discussions on small fluxes. We impose periodic boundary conditions for both $x,y$-directions in the torus geometry, for which the flux is quantized as $\phi=p/L_{x}$, with $p\in{\mathbb{Z}}$. On the other hand, the cylinder geometry has open (periodic) boundary conditions for the $x(y)$ direction. The flux is quantized as $\phi=p/\{(L_{x}-1)L_{y}\}$, for which (projective) mirror symmetry $x\to L_{x}-x$ is kept in the cylinder.
Since the Aharonov-Casher effect is a relativistic phenomenon, an experimentally achievable $\phi$ would be small. For example, when the lattice constant is 1 $\mathrm{\AA}$ and the electric field gradient $\mathcal{E}=200{\rm V/\mu m^{2}}$, the fictitious magnetic field is $\phi\sim 0.01$ and the magnetic length $l_{B}\sim 35.6{\rm V/\mu m}$. On the other hand, our DMRG analysis in Sec. III.2 reveals nontrivial behavior of the Néel order parameter even under a very small fictitious magnetic field. This implies that it would be possible to experimentally observe changes of the Néel order under the fictitious magnetic field. We note that an efficient method for achieving a moderate Aharonov-Casher effect has been proposed in a previous study [19]. Moreover, a fictitious magnetic field with a scale larger than the Aharonov-Casher effect could be applied to spin systems through artificial gauge fields in cold atomic systems [38, 39].
@@
In this paper, we first apply the spin-wave approximation, which is a standard method for describing the low-energy properties of quantum magnets. The Hamiltonian is approximated within quadratic terms in the boson operators and interactions between the magnons are neglected (see Appendix A). In the spin-wave approximation, both torus and cylinder geometries are used depending on physical quantities considered and ranges of the fictitious magnetic flux. Next, we perform the DMRG calculation in the cylinder geometry to obtain more accurate results. The quantum fluctuations arising from the magnon interaction neglected in the spin-wave approximation are considered in the DMRG calculation. The aspect ratio is chosen as $(L_{x},L_{y})=(2L,L)$, with the system size $L=6,8,\text{and}\,10$. The bond dimension is $\chi=800-3200$, and the truncation error is $\epsilon<10^{-5}$. [25, 24, 26].
@@ III Calculation Results
In this section, we show the calculation results of the spin-wave approximation and the DMRG at zero temperature. They give qualitatively consistent results, but there is significant quantitative difference. To have insight into the results of DMRG calculation , we give a brief discussion based on the magnetic catalysis.
@@ III.1 Spin-wave approximation
In the spin-wave approximation, we suppose that the classical ground state at $\phi=0$ is the Néel ordered state in the spin-$z$axis. Under a fictitious magnetic field, the system has U(1)×${\times\mathbb{Z}}_{2}$ symmetry, where U(1) symmetry breaking corresponds to Néel order in the $xy$-plane and ${\mathbb{Z}}_{2}$ symmetry breaking to Néel order along the $z$-axis. Therefore, the Néel order parameter along the $z$-axis is enhanced if ${\mathbb{Z}}_{2}$ symmetry breaking is more strongly favored in rhe presence of a fictitious magnetic field. On the other hand, the staggered moment will be aligned in the $xy$-plane if U(1) symmetry is spontaneously broken by the fictitious magnetic field, which means that the $z$-axis Néel order will be suppressed. In the latter case, the $z$-axis Néel ansatz is inappropriate and the spin-wave approximation based on this ansatz state would not work. More precisely, the energy spectrum of magnons will have imaginary components when the spin-wave approximation becomes invalid. A failure of the spin-wave approximation implies instability of the Néel ordered ansatz, and a different order may be realized in the ground state.
We first calculate the energy spectrum of the magnons and investigate how the magnon excitations depend on the fictitious magnetic flux $\phi$.
@@
Figure 2 shows the energy spectrum calculated by the spin-wave approximation for various magnetic fluxes $0\leq\phi\leq 1$. The magnon spectrum shows a Hofstadter butterfly as in a free fermion system, as demonstrated in previous studies [40, 21, 22]. Note that the calculated magnon energy is always real (nonimaginary) for all $\phi$, which implies that the $z$-axis Néel ground state is stable under the fictitious magnetic field. If the $z$-axis Néel state were unstable under some fluxes, the magnon energy would take complex values. One can see that the magnon excitation acquires an energy gap under the magnetic field $\phi\neq 0$. The magnon excitation gap $\Delta E$ (the smallest magnon excitation energy) behaves as $\Delta E\propto\sqrt{\phi}$ for small magnetic fluxes $0<\phi\lesssim 0.01$, as shown in Figs. 3(a) and (b), which is consistent with the long-wavelength behavior described by the Klein-Gordon equation [22].
Figure 3: Magnon energy gap for small values of the fictitious magnetic field (a) in the linear scale and (b) in the logarithmic scale. The system size is $L_{x}=L_{y}=100$ sites with the cylinder geometry. The dashed curve in (b) is $\Delta E\sim\sqrt{\phi}$. Small deviation from the exact $\Delta E\sim\sqrt{\phi}$ behavior is considered a finite-size effect.
Figure 4: (a) Néel order parameter $m_{s}$ calculated by the spin-wave approximation in the cylinder geometry for the system size $L_{x}=100,L_{y}=50$. The inset is the difference $\Delta m_{s}$. (b) Logarithmic scale plot of $\Delta m_{s}$. The dashed line represents $\Delta m_{s}\propto\sqrt{\phi}$.
The $\Delta E\propto\sqrt{\phi}$ dependence can be understood by a simple dimensional analysis of the linear spectra in the momentum space $\Delta E\propto k$ at $\phi=0$. Indeed, $\Delta E\propto k\propto 1/L$ for the system size $L$, and one can just replace $L$ with the magnetic length $l_{B}=1/\sqrt{\phi}$ to obtain $\Delta E\propto 1/l_{B}=\sqrt{\phi}$. Note that the long-wave length limit of the magnon spectrum is effectively described by the Klein-Gordon equation and there is no zero-mode ($\Delta E=0$) in sharp contrast with Dirac fermions at charge neutrality. The non-zero magnon gap implies that any excitations from the ground state are gapped since the magnons are elementary excitations in the Heisenberg model. The correlation length will be given by $\xi\sim v/\Delta E\propto l_{B}$, where $v$ is the speed of the magnon excitation, and the system is governed by this length scale.
@@
where we use use the cylinder geometry since it allows small fluxes. Here $N$ is the number of sites in the summation and it is restricted to the $N=L\times L$ sites in the middle of the system for the cylinder geometry with $(L_{x},L_{y})=(2L,L)$ to suppress contributions from the open boundary. Note that the spin-wave approximation works quite well for the Heisenberg model at $\phi=0$ and it gives $m_{s}\simeq 0.303$ in the thermodynamic limit which is very close to $m_{s}\simeq 0.306-0.309$ obtained by the extensive numerical calculations [41, 42, 43, 44, 45, 46]. Figure 4(a) shows the Néel order parameter $m_{s}$ under the fictitious magnetic flux $\phi$. We find that the Néel order parameter is enhanced from the zero-flux value $m_{s}(0)\simeq 0.303$, and it reaches $m_{s}\sim 0.41$ at large fluxes. Once the enhanced $m_{s}$ has been obtained, it can be naturally understood from the gapped magnon spectrum in analogy to an Ising-like system at $\phi=0$, where magnons are gapped, and correspondingly, $m_{s}$ is large. Since zero-point motion of the magnons leads to the suppression of $m_{s}$ from the classical value $m_{\rm max}=0.5$ within the spin-wave approximation, a magnon gap would generally imply increase of $m_{s}$. At the same time, however, the enhanced Néel order may be non-trivial, once one notices that the Hamiltonian in Eq.(1) partly includes a non-uniform Dzyaloshinskii-Moriya interaction in the spin $xy$-plane when $e^{i\theta_{ij}}$ is simply Taylor expanded up to the lowest order in $\phi$. A uniform $xy$-Dyzloshinskii-Moriya interaction is known to destabilize the $z$-axis Néel order, and one might naively expect that the non-uniform Dyzloshinskii-Moriya interaction has a similar influence. The enhancement of the Néel order in our calculation suggests that it is important to fully consider the Landau quantization beyond the naive Taylor expansion.